OpenStudy (anonymous):

Does (x^3 - x^-3)^2 = -2 ??

zepdrix (zepdrix):

\[\Large (x^3-x^{-3})^2\qquad=\qquad \left(x^3-\frac{1}{x^3}\right)^2\]Getting a common denominator in the brackets gives us,\[\Large \left(\frac{x^6-1}{x^3}\right)^2\] No it doesn't simplify down to -2 if that's what you were asking :)

OpenStudy (anonymous):

Okay, cool. Thanks :)